OpenStudy (anonymous):

(These questions are the ones i got wrong on my test I need someone to help me answer them please.) What are the solutions of the system? Using a graphing calculator. y=x^2+4.6x+4.7 y=3.3x-3.7 A. (-3.47,3.48) and (2.17, -15.17) C. (-3.47, -15.17) and (2.17,3.48) D. No solution

OpenStudy (turingtest):

you can use a graphing calculator?

OpenStudy (anonymous):

yes but i dont have one and the ones on the internet dont help

OpenStudy (turingtest):

if you have access to the internet you can use wolfram alpha: http://www.wolframalpha.com/ http://www.wolframalpha.com/input/?i=plot%20y%3Dx%5E2%2B4.6x%2B4.7%2C%20y%3D3.3x-3.7%20&t=crmtb01

OpenStudy (anonymous):

like i dont know how to put it in the calculator

OpenStudy (turingtest):

that depends on the calculator I suppose I wouldn't know really how to use them, I use wolfram usually notice all I did was put "plot" with the equations separated by a comma and you can see them drawn out do they intersect?

OpenStudy (anonymous):

ok i got that part but how do i get the answer

OpenStudy (turingtest):

what does it mean graphically to solve the system? -it means finding the intersection of the two graphs no calculator is really required x^2+4.6x+4.7=3.3x-3.7 x^2+1.3x+8.4=0 now find the discriminant of this quadratic; if it is negative that means the equation has two complex roots, so there are no real intersections...

OpenStudy (turingtest):

if there were real solutions to this equation you would plug in the answers you get for x into each equation and get a set of ordered pairs

OpenStudy (anonymous):

well I think it is no solution but im not really sure and thats what i was thinking bout the calculator it makes no since

OpenStudy (turingtest):

\[ax^2+bx+c=0\implies x={-b\pm\sqrt{b^2-4ac}\over2a}\]the discriminant is\[b^2-4ac\]when it is >0 there are two distinct real roots when it =0 there is a double real root when it is <0 it is two complex roots

OpenStudy (turingtest):

if you can graph it the image is enough to see that they do not intersect since the vertex of the quadratic does not go below the line

OpenStudy (anonymous):

so is the answer no solution

OpenStudy (turingtest):

right

OpenStudy (anonymous):

ok thank you so much you really helped me

OpenStudy (turingtest):

no problem :)

OpenStudy (anonymous):

maybe you can help witht he next one im about to post

OpenStudy (anonymous):

*the

OpenStudy (turingtest):

maybe, but I'm doing a few things at once please post separately each problem, many others can help you I promise